OpenStudy (anonymous):

Suppose that f(X)=x^2-3x-18 and g(x)=x^2+2x-3 a. Find the zeros of (f+g)(x) b. Find the zeros of (f-g)(x) c. Find the zeros of (f times g)(x)

OpenStudy (anonymous):

\[\left( f+g \right)\left( x \right)=f \left( x \right)+g \left( x \right)\] simplify and find solve for x